2. A convention of sumo wrestlers is held at a hotel. The weight of sumo wrestlers can be approximated by a normal with mean 540 pounds and standard deviation of 45 pounds. The hotel elevator can accommodate 9 wrestlers at a time. Suppose that a simple random sample of 9 wrestlers enters the elevator. The elevator will fail if the total weight of the occupants exceeds 5000 pounds. What is the chance that the elevator will fail to operate?
1
Expert's answer
2019-10-29T10:07:44-0400

If the total weight of the 9 wrestlers exceeds 5000 pounds, then the mean weight "\bar{x}" of the 9 wrestlers must exceed "5000\/9" pounds.

Change the criterion for failing to operate from "total weight exceeding 5000 pounds" to "mean weight for 9 people exceeding "5000\/9" pounds".

Let "X=" the mean weight in pounds: "X\sim N(\mu,\sigma^2)"

Then


"Z={X-\mu \over \sigma}\sim N(0,1)"

Given that "\mu=540" pounds, "\sigma=45/\sqrt{9}" pounds.


"P(X>{5000 \over 9})=1-P(X\leq{5000 \over 9})="

"=1-P(Z\leq{{5000 \over 9}-540 \over 15})\approx1-0.850141\approx0.1499"

The probability that the elevator will fail to operate is "0.1499."